# EDEN: Exploring Disks Embedded in N-body simulations of Milky-Way-mass halos from Symphony

### v1

@@ 5.4 Effects of numerical resolution
As mentioned in Section 4.3, the quantitative results in this work, such as the subhalo counts ratio, SPMF, and radial distributions, may be subject to change under different numerical resolutions. We show in Appendix D that, indeed, resolution matters, as well as the specific choice of the subhalo finder. We re-simulate five SymphonyMilkyWay hosts at 8×8$\times$8 × higher resolution. Higher resolution leads to a higher SPMF that boosts subhalo counts at all mass scales. The boost in subhalo counts due to higher numerical resolution is comparable to the magnitude of disk effects on subhalo abundances at the same fiducial resolution, further highlighting that the results of disk effects are sensitive to numerical resolution.
Therefore, we advocate for future work quoting quantitative comparisons to EDEN to be conducted at similar numerical resolution and to apply the same halo finder (Symfind). As we have checked, the subhalo counts ratios shown in Fig. 5 are more consistent between Symfind and Rockstar than the SPMF (Fig. 11) which is more sensitive to the specific halo finder choice. Qualitative comparisons between different simulations using different halo finders should prioritize subhalo abundance ratio ($N_{\rm Disk}/N_{\rm DMO}$) comparisons rather than absolute subhalo abundance (SPMF).
@@ 5.5 The Milky Way’s disk in a cosmological context
The most important finding in this work is that subhalo abundance suppression in MW-mass halos is sensitive to the disk-to-halo mass ratio, and thus the MW (as well as M31), with a larger-than-average disk, likely suppresses subhalos more efficiently (Fig. 5). This has profound implications for understanding the galaxy–halo connection of low-mass satellites around MW-mass hosts external to the MW. If one tries to constrain galaxy–halo connection on a broad set of MW-mass objects in the Local Universe from Surveys such as SAGA (Geha et al., 2017; Mao et al., 2021) and ELVES (Carlsten et al., 2022), the overall effect of subhalo abundance suppression across their host samples would be weaker than in the MW itself, due to the rareness ($\sim 1\sigma$ outlier) of a MW-like heavy stellar disk.
For instance, the MW is a $\gtrsim 1\sigma$ larger-than-average host compared to the 101 hosts in SAGA (see Fig. 2 in Wang et al. 2024), consistent with EDEN hosts shown in Figs. 1 and 5. For the majority of SAGA-host, that coincides with $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$, disk effects of $\sim 10\%$ within $R_{\rm vir,host}$ and $\sim 20\%$ within 100 kpc are expected for subhalos according to EDEN (Fig. 5). In those cases, constraints on the low-mass galaxy–halo connection based on average number densities of satellites within $R_{\rm vir,host}$ is rather robust given the expected $\sim 10\%$ disk effects, but radial trends could change significantly especially within the central 100 kpc due to stronger suppression at smaller radii ($\gtrsim 20$, Fig. 7). This indicates that EDEN might mitigate the limitation of UM-SAGA, which predicts fewer quenched satellites than SAGA observations within 100 kpc (Wang et al., 2024), by accounting for the additional tidal forces from the host galactic disk.
@@
In the near future, we plan to extend the EDEN framework to the new ‘MW-est’ zoom-in simulations (Buch et al., 2024), a suite of 20 MW-like zoom-in simulations that have similar halo masses as SymphonyMilkyWay hosts but also have tailored Gaia-Sausage Enceladus (Belokurov et al., 2018; Helmi et al., 2018) and LMC (van der Marel et al., 2002; Besla et al., 2007; Kallivayalil et al., 2013; Vasiliev, 2023) mergers. We also plan to use EDEN as the backbone to model Local Group low-mass galaxies (Weisz et al., 2014; Savino et al., 2023) along with SAGA satellites over the entire observable galaxy stellar mass range ($M_{\ast}>10^{2}\mathrm{M_{\astrosun}}$), setting the MW and M31 satellites in a broader cosmological context and constraining the process of cosmic reionization.
@@ 6 Conclusions
A novel feature of the work is that the analytic disk potentials grow self-consistently with their equivalent stellar masses predicted by the empirical galaxy–halo connection model UniverseMachine (Behroozi et al., 2019) (Section 2). This ensures that the broad range of disk masses and disk formation histories are well-sampled in a correlated fashion with halo mass and halo assembly. Our simulation suite focuses on $M_{\rm vir}\sim 10^{12.1}\mathrm{M_{\astrosun}}$ halos and covers heavy disk systems like our MW (Fig. 2). We use the particle-tracking subhalo finder Symfind (Mansfield et al., 2023) to track subhalos down to $M_{\rm peak}\geqslant 300m_{\rm DM}=1.2\times 10^{8}\mathrm{M_{\astrosun}}$ and quantify how the subhalo populations change in the presence of a central disk potential. We have also re-simulated nine embedded-disk halos with ×$\times 2.5$ heavier disks (EDEN $M_{\ast,\rm Disk}\times 2.5$) to explore the effect of disk mass while fixing the disk growth history. Our main findings are:
@@
The most important findings of this work are the first two points mentioned above. It is clear from Figs. 1 and 5 that the MW has an atypically large disk for an average $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ halo. Previous simulation work (D’Onghia et al., 2010; Garrison-Kimmel et al., 2017; Kelley et al., 2019) with embedded disk potentials implementing MW-like heavy stellar disks ($M_{\ast}\gtrsim 6\times 10^{10}\mathrm{M_{\astrosun}}$) are potentially biased towards heavy disk systems and overestimates the average subhalo abundance suppression intensities in MW-mass halos. Subhalo abundance suppression is less significant in a broader range of MW-mass halos in the Universe, and suppression ratios obtained in previous models using heavy disk should not be naively generalized to the broader range of MW-mass hosts at $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$. The disk effects on subhalo abundances at other host halo mass scales significantly smaller or larger than $M_{\rm vir,host}\sim 10^{12}\mathrm{M_{\astrosun}}$ are likely even less important due to their smaller $M_{\ast,\rm Disk}/M_{\rm vir,host}$ (Wechsler & Tinker, 2018).
@@ Acknowledgments
We plan to release EDEN halo catalogs and merger trees upon publication of this manuscript.
We thank James Bullock, Benedikt Diemer, Pratik Gandhi, Yao-Yuan Mao, Volker Springel, Frank van den Bosch, Andrew Wetzel and the GFC group at KIPAC/Stanford for helpful discussion and insightful comments during the preparation of this draft. The FIRE-2 cosmological zoom-in simulations of galaxy formation are part of the Feedback In Realistic Environments (FIRE) project, generated using the Gizmo code (Hopkins, 2015) and the FIRE-2 physics model (Hopkins et al., 2018).
@@ Appendix C Host trajectories and subhalo infall times
As discussed in Section 5.3, the simplistic treatment of fixing the disk orientation may impact the halo spin and its 3D position through exerting a torque on the dark matter particles. In this Appendix we show how well the sink particle is anchored to the halo center (potential minimum) and how much the added disk potential changes the spatial trajectory of the host halo.
@@
However in the bottom panel, we observe that the distance between the halo center of the MW host halo in the DMO run starts to deviate from the Disk run since $z=3$ and reaches a median distance of 12 kpc at $z=0$. Five halos end up more than 30 kpc away from their location in the DMO run once the embedded disk potential is added, which is $\gtrsim 10\%$ of their halo $R_{\rm vir}$. The Phat-ELVIS team also kindly provided us trajectory data for one of their ten halos (private comm. James Bullock, Hyun-su Kim). It also shows a 27 kpc offset between their DMO and embedded Disk halos at $z=0$. Therefore, this systematic offset is a generic caveat for embedded disk simulations with fixed disk orientation.
@@ Appendix D Resolution tests and comparison between halo finders
As mentioned in Section 5.3, the quantitative results of subhalo abundances (i.e., the subhalo counts ratio) might be sensitive to the numerical resolution of the simulation. We show in Fig. 11 how numerical resolution and the specific choice of the halo finder affects the subhalo peak mass function quantitatively.
In the left panel, we compare SPMF in Symfind halo catalogs for Symphony and EDEN fiducial halos at their fiducial resolution (8K) and 8×8$\times$8 × higher resolution (16K). Comparing the Symphony 8K and 16K curves, more subhalos will be tracked and survive till $z=0$ at all subhalo mass scales with $M_{\rm peak}>300m_{\rm DM}=1.2\times 10^{8}\mathrm{M_{\astrosun}}$. The change in SPMF due to the change in resolution is comparable to the disk effects on subhalo abundance at the fiducial 8K resolution. This indicates that the EDEN SPMF and subhalo counts ratio are not fully converged under the current resolution, and that the quantitative arguments in this work on subhalo abundance suppression is sensitive to numerical resolution. We emphasize that the qualitative conclusion of subhalo abundance suppression happening at all subhalo mass scales are unchanged with increasing resolution. We plan to re-simulate more EDEN and Symphony hosts in the future at 16K resolution.
@@ Appendix E Comparing to the N18 random forest model
To connect our model predictions with latest observational constraints on subhalo abundances, we apply the random-forest (RF) model from Nadler et al. (2018) (N18 for short) to the SymphonyMilkyWay DMO halos. The RF model was trained on the hydro–DMO pairs of the two halos introduced in Garrison-Kimmel et al. (2017), which were run using the FIRE-2 hydrodynamic simulation code (Hopkins et al., 2018). The model uses 5 subhalo properties (pericenter distance, pericenter scale factor, infall mass, infall scale factor, and infall maximum circular velocity) from the DMO run to predict the probability that their counterparts in the corresponding hydrodynamic simulations are lost due to tidal stripping. Since Garrison-Kimmel et al. (2017) used Rockstar halo catalogs, we also apply the N18 RF model to the SymphonyMilkyWay Rocsktar catalogs. Given the large fraction of subhalos numerically lost track by Rockstar (Fig. 11), we caution the reader to treat this subhalo abundance suppression comparison of the RF model and actual disk potentials as a qualitatively illustration.
@@
The results of this model applied to Symphony DMO hosts are shown in Fig. 13. The RF model applied on the full suite of 45 MW-mass halos are consistent with the SPMFs of the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ within $100$ kpc, except for subhalos at $M_{\rm peak}\sim 6\times 10^{9}\mathrm{M_{\astrosun}}$. The difference at this subhalo mass scale is mainly due to switching to Rockstar catalogs for EDEN $M_{\ast,\rm Disk}\times 2.5$, which may have introduced ‘merger tree’ stitching errors that assigned lower mass progenitors to higher mass descendants after pericenter passage (Mansfield et al., 2023). Nonetheless, the RF agrees well with EDEN $M_{\ast,\rm Disk}\times 2.5$ apart from this mass scale and loses significantly more lower-mass subhalos ($M_{\rm peak}\lesssim 10^{9}\mathrm{M_{\astrosun}}$) than EDEN fiducial. This agreement with the high-disk-mass subsample in EDEN is expected as the RF model was trained on the two Garrison-Kimmel et al. (2017) halos with large MW-like disk masses that leads to strong subhalo abundance suppression (Fig. 5 top panel).

### v3

@@ 5.4 Effects of numerical resolution
As mentioned in Section 4.3, the quantitative results in this work, such as the subhalo count ratio, SPMF, and radial distributions, may be subject to change under different numerical resolutions. We show in Appendix D that, indeed, resolution matters, as well as the specific choice of the subhalo finder. We re-simulate five SymphonyMilkyWay hosts at 8×8$\times$8 × higher resolution. Higher resolution leads to a higher SPMF that boosts subhalo counts at all mass scales. The boost in subhalo counts due to higher numerical resolution in the dark-matter-only case is comparable to the magnitude of disk stripping effects on subhalo abundances at the same fiducial resolution, further highlighting that the results of disk effects are sensitive to numerical resolution.
The particle-tracking-based Symfind halo finder performs much better at tracking heavily stripped subhalos than the phase space overdensity-based halo finder Rockstar (see Section 4.2 and Fig. 5 in Mansfield et al. 2024). The key takeaway is that Rockstar misses a sizable fraction of $\lesssim 10^{5}$ peak-particle-count subhalos (Fig. 10, Mansfield et al., 2024), and the majority of the subhalos that Rockstar misses are heavily stripped. As a result, Rockstar yields false numerical convergence for the abundance of subhalos with $\sim 100$ to $10^{5}$ peak particles. The cause of this is rooted in the nature of the phase-space overdensity becoming less prominent as a subhalo starts being stripped; these overdensities thus become much harder to identify in Rockstar than the most-bound particles used in Symfind. This problem is likely even more severe for physical-overdensity-based halo finders such as Subfind (Springel et al., 2001; Dolag et al., 2009), where only spatial density is used to identify subhalos (see also the discussion in Appendix H of Mansfield et al. 2024).
For heavily stripped subhalos in our fiducial-resolution EDEN simulations, we show that Symfind recovers more $\gtrsim 10^{4}$ peak particle count subhalos than Rockstar when resolution is increased (Appendix D); this result is consistent with Fig. 10 of Mansfield et al. (2024). This is a regime where previous studies have long believed that phase-space and spatial overdensity halo finders are complete and converged (e.g., Power et al., 2003; Diemand et al., 2004; Reed et al., 2005; Springel et al., 2008; Navarro et al., 2010; Gao et al., 2012; Ludlow et al., 2019a; Mansfield & Avestruz, 2021; Nadler et al., 2023b), whereas they are in fact underestimating subhalo abundances compared to idealized simulations (Peñarrubia et al., 2010; van den Bosch et al., 2018; van den Bosch & Ogiya, 2018; Errani & Peñarrubia, 2020; Errani & Navarro, 2021).
Nevertheless, we emphasize that the EDEN results—and particularly the $N_{\rm Disk}/N_{\rm DMO}$ predictions—are more robust than previous studies given our current numerical resolution. This follows as we focus on relative subhalo abundances in disk versus DMO runs rather than absolute abundances, paired with particle-tracking subhalo finder Symfind. As we have checked, the subhalo count ratios shown in Fig. 5 are more consistent between Symfind and Rockstar than the SPMF (Fig. 11) which is more sensitive to the specific halo finder choice. We suggest that future work which includes quantitative comparisons to EDEN results should control for numerical resolution and subhalo finder differences. Qualitative comparisons between different simulations using different halo finders should prioritize subhalo abundance ratio ($N_{\rm Disk}/N_{\rm DMO}$) comparisons rather than absolute subhalo abundance (SPMF).
@@ 5.5 The Milky Way’s disk in a cosmological context
The most important finding in this work is that subhalo abundance suppression in MW-mass halos is sensitive to the disk-to-halo mass ratio, and thus the MW (as well as M31), with a larger-than-average disk, likely suppresses subhalos more efficiently (Fig. 5). This has profound implications for understanding the galaxy–halo connection of low-mass satellites around MW-mass hosts external to the MW. If one tries to constrain the galaxy–halo connection on a broad set of MW-mass objects in the Local Universe from Surveys such as SAGA (Geha et al., 2017; Mao et al., 2021, 2024; Geha et al., 2024; Wang et al., 2024) and ELVES (Carlsten et al., 2022), the overall effect of subhalo abundance suppression across their host samples would be weaker than in the MW itself, due to the rareness ($\sim 1\sigma$ outlier) of a MW-like heavy stellar disk.
For instance, the MW is a $\gtrsim 1\sigma$ larger-than-average host compared to the 101 hosts in SAGA (see Fig. 2 in Wang et al. 2024), consistent with EDEN hosts shown in Figs. 1 and 5. For the majority of SAGA hosts, which coincide with $M_{\ast,\rm Disk}/M_{\rm vir,host}\lesssim 0.02$, disk effects of $\sim 10\%$ within $R_{\rm vir,host}$ and $\sim 20\%$ within 100 kpc are expected for subhalos according to EDEN (Fig. 5). In those cases, constraints on the low-mass galaxy–halo connection based on average number densities of satellites within $R_{\rm vir,host}$ are rather robust given the expected $\sim 10\%$ disk effects, but radial trends could change significantly especially within the central 100 kpc due to stronger suppression at smaller radii ($\gtrsim 20$, Fig. 7). This indicates that EDEN might mitigate the limitation of UM-SAGA, which predicts fewer quenched satellites than SAGA observations within 100 kpc (Wang et al., 2024), by accounting for the additional tidal forces from the host galactic disk.
@@
In this context, it is interesting to compare the $M_{\ast,\rm Disk}/M_{\rm vir,host}$ ratios and satellite counts in the MW and M31, which have 5 and 8 satellites with $M_{\ast}>10^{7}\mathrm{M_{\astrosun}}$, respectively (McConnachie, 2012; Drlica-Wagner et al., 2020). According to Fig. 5, M31 has a higher $M_{\ast,\rm Disk}/M_{\rm vir,host}$ than the MW and should thus have a lower $N_{\rm disk}/N_{\rm DMO}$ due to stronger tidal stripping. Some results suggest that the MW ($1.08^{+0.14}_{-0.20}\times 10^{12}\mathrm{M_{\astrosun}}$ Cautun et al., 2020) and M31 ($1.4^{+0.4}_{-0.4}\times 10^{12}\mathrm{M_{\astrosun}}$ Watkins et al., 2010) have similar halo masses, although the uncertainties allow for MW/M311:2:similar-toabsent12$\sim$ 1:2∼ 1 : 2 virial mass ratios. [^9] The Local Group “timing argument” also allows for a significantly larger M31 virial mass, e.g., Benisty 2024. Nevertheless, given the similar mean-inferred halo masses, we should also consider secondary halo bias and intrinsic host-to-host scatter in subhalo counts at fixed halo mass, which is embedded in the scatter of $N_{\rm Disk}/N_{\rm DMO}$ in Fig. 5. As shown in the SAGA Survey Paper III (Mao et al., 2024), the observed satellite count in MW-mass hosts is predominantly correlated with the stellar mass of their most massive satellite (Fig. 15). Recent observations (D’Souza & Bell, 2021) indicate that M31’s most massive satellite M32 could have a stellar mass as large as $2.5\times 10^{10}\mathrm{M_{\astrosun}}$, making it almost an order-of-magnitude heavier than the LMC ($2.7\times 10^{9}\mathrm{M_{\astrosun}}$, van der Marel et al., 2002). This could mean that M31 had a much higher $N_{\rm DMO}$ to start with, even if its virial mass is comparable to the MW. In this scenario, when M31’s subhalos go through much stronger tidal stripping due to its heavier disk, M31 still ends up having a higher $N_{\rm disk}$ than the MW while maintaining a lower $N_{\rm Disk}/N_{\rm DMO}$.
Generalizing the disk impact from MW-mass hosts to other host mass scales, we expect subhalo abundance suppression due to disk tidal stripping to be weaker in other halo mass scales given that the stellar-to-halo mass ratio is higher around $M_{\rm vir,host}\sim 10^{12}\mathrm{M_{\astrosun}}$ MW-mass halos (e.g., Behroozi et al., 2013a, 2019; Nadler et al., 2023b; Wang et al., 2024). However, subhalo abundance in group-mass halos $M_{\rm vir,host}\sim 10^{13}\mathrm{M{\astrosun}}$ might be affected by MW-mass progenitors whose pre-processed (Wetzel et al., 2015) subhalo abundance might carry over. This is an interesting topic to be explored in the future, but we speculate that it might be highly dependent on the specific assembly history of the group host, i.e., at what redshift those MW-sized satellites fell in, and whether their galaxies grew massive enough to significantly affect their satellite population at their accretion redshift. Dedicated zoom-in simulations embedding galaxies into both the Symphony group-mass hosts and their MW-mass infall progenitors would be required to answer these questions.
In the near future, we plan to extend the EDEN framework to the new ‘MW-est’ zoom-in simulations (Buch et al., 2024), a suite of 20 MW-like zoom-in simulations that have similar halo masses as SymphonyMilkyWay hosts but also have tailored Gaia-Sausage Enceladus (Belokurov et al., 2018; Helmi et al., 2018) and LMC (van der Marel et al., 2002; Besla et al., 2007; Kallivayalil et al., 2013; Vasiliev, 2023) mergers.
@@ 6 Conclusions
A novel feature of the work is that the analytic disk potentials grow self-consistently with their equivalent stellar masses predicted by the empirical galaxy–halo connection model UniverseMachine (Behroozi et al., 2019) (Section 2). This ensures that the broad range of disk masses and disk formation histories are well-sampled in a correlated fashion with halo mass and halo assembly. Our simulation suite focuses on $M_{\rm vir}\sim 10^{12.1}\mathrm{M_{\astrosun}}$ halos and covers heavy disk systems like our MW (Fig. 2). We use the particle-tracking subhalo finder Symfind (Mansfield et al., 2024) to track subhalos down to $M_{\rm peak}\geqslant 300m_{\rm DM}=1.2\times 10^{8}\mathrm{M_{\astrosun}}$ and quantify how the subhalo populations change in the presence of a central disk potential. We have also re-simulated nine embedded-disk halos with ×$\times 2.5$ heavier disks (EDEN $M_{\ast,\rm Disk}\times 2.5$) to explore the effect of disk mass while fixing the disk growth history. Our main findings are:
@@
  5. 5.
    We advocate that comparisons between different simulations should prioritize comparing subhalo abundance ratio ($N_{\rm Disk}/N_{\rm DMO}$) rather than subhalo mass functions to mitigate subhalo finder systematics.
The most important findings of this work are the first two points mentioned above. It is clear from Figs. 1 and 5 that the MW has an atypically large disk for an average $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$ halo. Previous simulation work (D’Onghia et al., 2010; Garrison-Kimmel et al., 2017; Kelley et al., 2019) with embedded disk potentials implementing MW-like heavy stellar disks ($M_{\ast}\gtrsim 6\times 10^{10}\mathrm{M_{\astrosun}}$) is potentially biased towards heavy disk systems and overestimates the average subhalo abundance suppression intensities in MW-mass halos. Subhalo abundance suppression is less significant in a broader range of MW-mass halos in the Universe, and suppression ratios obtained in previous models using heavy disks should not be naively generalized to the broader range of MW-mass hosts at $M_{\rm vir}\sim 10^{12}\mathrm{M_{\astrosun}}$. The disk effects on subhalo abundances at other host halo mass scales significantly smaller or larger than $M_{\rm vir,host}\sim 10^{12}\mathrm{M_{\astrosun}}$ are likely even less important due to their smaller $M_{\ast,\rm Disk}/M_{\rm vir,host}$ (Wechsler & Tinker, 2018).
@@ Acknowledgments
We thank James Bullock, Benedikt Diemer, Pratik Gandhi, Yao-Yuan Mao, Volker Springel, Frank van den Bosch, Andrew Wetzel and the GFC group at KIPAC/Stanford for helpful discussions and insightful comments during the preparation of this draft. The FIRE-2 cosmological zoom-in simulations of galaxy formation are part of the Feedback In Realistic Environments (FIRE) project, generated using the Gizmo code (Hopkins, 2015) and the FIRE-2 physics model (Hopkins et al., 2018).
@@ Appendix C Host trajectories and subhalo infall times
As discussed in Section 5.3, the simplistic treatment of fixing the disk orientation may impact the halo spin and its 3D position by exerting a torque on the dark matter particles. In this Appendix, we show how well the sink particle is anchored to the halo center (potential minimum) and how much the added disk potential changes the spatial trajectory of the host halo.
@@
However, in the bottom panel, we observe that the distance between the halo center of the MW host halo in the DMO run starts to deviate from the Disk run since $z=3$ and reaches a median distance of 12 kpc at $z=0$. Five halos end up more than 30 kpc away from their location in the DMO run once the embedded disk potential is added, which is $\gtrsim 10\%$ of their halo $R_{\rm vir}$. The Phat-ELVIS team also kindly provided us trajectory data for one of their ten halos (private comm. James Bullock, Hyun-su Kim). It also shows a 27 kpc offset between their DMO and embedded Disk halos at $z=0$. Therefore, this systematic offset is a generic caveat for embedded disk simulations with fixed disk orientation.
@@ Appendix D Resolution tests and comparison between halo finders
As mentioned in Section 5.4, the quantitative results of subhalo abundances (i.e., the subhalo count ratio) might be sensitive to the numerical resolution of the simulation. We show in Fig. 11 how numerical resolution and the specific choice of the halo finder affect the subhalo peak mass function quantitatively.
In the left panel, we compare SPMF in Symfind halo catalogs for Symphony and EDEN fiducial halos at their fiducial resolution (8K) and 8×8$\times$8 × higher resolution (16K). Comparing the Symphony 8K and 16K curves, more subhalos will be tracked and survive till $z=0$ at all subhalo mass scales with $M_{\rm peak}>300m_{\rm DM}=1.2\times 10^{8}\mathrm{M_{\astrosun}}$. The change in SPMF due to the change in resolution is comparable to the disk effects on subhalo abundance at the fiducial 8K resolution. This indicates that the EDEN SPMF and subhalo count ratio are not fully converged under the current resolution, and that the quantitative arguments in this work on subhalo abundance suppression are sensitive to numerical resolution. We emphasize that the qualitative conclusion of subhalo abundance suppression happening at all subhalo mass scales is unchanged with increasing resolution. We plan to re-simulate more EDEN and Symphony hosts in the future at 16K resolution.
@@ Appendix E Comparing to the N18 random forest model
To connect our model predictions to latest observational constraints on subhalo abundances, we apply the random-forest (RF) model from Nadler et al. (2018) (N18 for short) to the SymphonyMilkyWay DMO halos. The RF model was trained on the hydro–DMO pairs of the two halos introduced in Garrison-Kimmel et al. (2017), which were run using the FIRE-2 hydrodynamic simulation code (Hopkins et al., 2018). The model uses 5 subhalo properties (pericenter distance, pericenter scale factor, infall mass, infall scale factor, and infall maximum circular velocity) from the DMO run to predict the probability that their counterparts in the corresponding hydrodynamic simulations are lost due to tidal stripping. Since Garrison-Kimmel et al. (2017) used Rockstar halo catalogs, we also apply the N18 RF model to the SymphonyMilkyWay Rocsktar catalogs. Given the large fraction of subhalos numerically lost track by Rockstar (Fig. 11), we caution the reader to treat this subhalo abundance suppression comparison of the RF model and actual disk potentials as a qualitatively illustration.
@@
The results of this model applied to Symphony DMO hosts are shown in Fig. 13. The RF model applied on the full suite of 45 MW-mass halos is consistent with the SPMFs of the nine halos in EDEN $M_{\ast,\rm Disk}\times 2.5$ within $100$ kpc, except for subhalos at $M_{\rm peak}\sim 6\times 10^{9}\mathrm{M_{\astrosun}}$. The difference at this subhalo mass scale is mainly due to switching to Rockstar catalogs for EDEN $M_{\ast,\rm Disk}\times 2.5$, which may have introduced ‘merger tree’ stitching errors that assigned lower mass progenitors to higher mass descendants after pericenter passage (Mansfield et al., 2024). Nonetheless, the RF agrees well with EDEN $M_{\ast,\rm Disk}\times 2.5$ apart from this mass scale and loses significantly more lower-mass subhalos ($M_{\rm peak}\lesssim 10^{9}\mathrm{M_{\astrosun}}$) than EDEN fiducial. This agreement with the high-disk-mass subsample in EDEN is expected as the RF model was trained on the two Garrison-Kimmel et al. (2017) halos with large MW-like disk masses that lead to strong subhalo abundance suppression (Fig. 5 top panel).
Appendix F The impact of the gaseous disk
In the EDEN disk potential setup (Section 2.3), we have kept the ‘gaseous’ disk potential a fixed fraction of the total disk mass throughout $z=3$ to $z=0$ (Eq. 1). We assumed that the EDEN total disk mass follows the UM-predicted stellar mass and did not account for any additional gas mass in order to make fairer comparisons with previous embedded disk simulations (Garrison-Kimmel et al., 2017; Kelley et al., 2019). Garrison-Kimmel et al. (2017) has shown that adding an additional gas disk that follows the gaseous disk growth in their hydrodynamic counterpart does not affect subhalo abundance (see their Fig. 7). They suggested that this could be due to the redshift mismatch between when the gas disk constituted a significant fraction of the baryonic mass at $z\gtrsim 1$ (e.g., Zhang et al., 2021; Chowdhury et al., 2022) and the majority of subhalos that survive to $z=0$ falling in at $z\lesssim 1$ (Wetzel et al., 2015), diminishing the impact of the gaseous disk component.
To further verify this finding, we randomly select three hosts out of the nine high-mass subsample EDEN hosts in Section 2.4 and re-simulate their disk potentials with an additional gaseous disk setup. We first renormalize the masses of the stellar disk $\Phi_{\ast}(R,Z)$ and the stellar bulge $\Phi_{b}(r)$ in Eq. 1 such that they sum up to the total stellar mass $M_{\ast,\rm UM}(z)$ given by UM at every snapshot (Fig. 2). Next, we normalize the gaseous disk potential mass in a time-varying fashion with respect to the UM-predicted stellar masses based on cosmic baryon densities predicted by the empirical gas–halo connection model NeutralUniverseMachine (Guo et al., 2023). NeutralUniverseMachine is an empirical gas–halo connection model anchored on the star formation histories of the UM DR1 model (Behroozi et al., 2019) and matches the HI/$\mathrm{H_{2}}$-stellar mass functions, HI-halo mass function, molecular-to-atomic ratio, and cosmic HI/$\mathrm{H_{2}}$ densities from $z=6$ to $z=0$. For simplicity, we assume that the adjusted gas disk mass $M_{\rm gas}(z)$ follows a time-varying ratio with respect to the UM-predicted stellar mass $M_{\ast,\rm UM}(z)$:
$\frac{M_{\rm gas}(z)}{M_{\ast,\rm UM}(z)}=\frac{1.36\rho_{\rm HI}(z)+\rho_{\rm H_{2}}(z)}{\rho_{\ast}(z)},$ (F1)
where $\rho_{\rm HI}(z)$, $\rho_{\rm H_{2}}(z)$, and $\rho_{\ast}(z)$ are the cosmic atomic hydrogen, molecular hydrogen, and stellar mass densities predicted by NeutralUnvierseMachine (their Fig. 11), and the 1.36 factor accounts for elements heavier than Hydrogen.
This is a limiting model, and for simplicity does not account for the variations in gas fraction across different halo mass scales. Because of this, we are likely to be overestimating the impact of gas disks since a MW-mass halo will often have less gas content than the Guo et al. (2023) average. However, given the modeling uncertainty in the evolution of gas content, testing a slightly extreme model can provide an upper bound estimate on the subhalo abundance suppression levels without loss of generality. We show the fiducial UM-predicted total disk mass and the gas-adjusted total disk mass histories along with the stellar mass fraction of the adjusted disks according to Eq. F1 for the three selected EDEN hosts in Fig. 14.
Figure 14: Main panel: Total disk mass history of the fiducial UM-predicted stellar mass (solid) and the gas-adjusted disk masses that follow time-varying gas mass ratios according to Eq. F1 (dashed). The slight decreasing trend in parts of the stellar mass histories is a manifestation of UM accounting for massive stars’ deaths over time. Inset: The implied universal stellar mass fraction $f_{\star}$ is inferred from the NeutralUniverseMachine-predicted cosmic HI, $\rm{H_{2}}$, and stellar mass densities (see Fig. 11 in Guo et al. 2023.)
Figure 15: Subhalo count ratios within 100 kpc (top) and $R_{\rm vir,host}$ (bottom) for the three adjusted gas-disk EDEN re-simulations (crosses). The gas-adjusted disk re-simulations are connected to their fiducial EDEN counterparts using dashes lines. The $M_{\ast}\times 2.5$ high stellar-mass counterparts are also shown for reference (diamonds, connected to fiducial EDEN using solid lines). The three gas-adjusted re-simulated halos shown in the same color coding as their mass history curves in Fig. 14. The adjusted gas disk with higher gas masses at earlier redshift has a smaller impact on $N_{\rm Disk}/N_{\rm DMO}$ compared to the mass normalization of the stellar disk.
We show in Fig. 15 the subhalo count ratios of the three re-simulations of EDEN hosts with gas-adjusted disk potentials. We compare the subhalo abundance suppression ratio $N_{\rm Disk}/N_{\rm DMO}$ of the gas-adjusted re-simulations with their fiducial EDEN runs and their ×$\times 2.5$ stellar disk mass re-simulations. Two of the three hosts we re-simulate with the adjusted gas disk potential show reduced $N_{\rm Disk}/N_{\rm DMO}$ compared to their fiducial EDEN counterparts, while one of them shows increased $N_{\rm Disk}/N_{\rm DMO}$. We summarize that:
  • The overall effect of adding a gas disk is similar to moving along the fitted $N_{\rm Disk}/N_{\rm DMO}$ relation, with the $z=0$ baryonic mass being the primary driver of subhalo abundance suppression without needing to differentiate between how much mass is in stars versus gas.
  • There is scatter in the amount of subhalo suppression/boost along the fitted relation. We cannot rule out that this scatter has some underlying physical cause (i.e., additional contraction by the gas disk) which needs a larger sample to verify. However, we note that structure formation is chaotic in nature and the gas disk could trigger a chaotic response that significantly changes subhalo infall times. Because of this we urge caution in over-interpreting individual hosts.
  • Given that the stellar and gaseous disks evolve differently, one explanation to the gas disk insensitivity could be that most subhalos fall in at $z\lesssim 1$ when the stellar disk starts to dominate ($f_{\ast}>50\%$, Fig. 14).
In all, these findings are consistent with the findings in Garrison-Kimmel et al. (2017), where the gas disk is a less decisive factor of subhalo abundance suppression due to disk tidal stripping. This also makes our choice of a fixed-gas-ratio choice for the disk model more justifiable in Section 2.3. Although the gas disk appears to be less decisive, our simplistic gas disk model and limited sample size do not rule out a more detailed study in the impact of the gaseous disk from being fruitful. In the future, we could explore the impact of the gas disk with higher-resolution simulations on a larger sample of hosts to further nail down systematics.